### v2

@@ Appendix B Dutch Auctions (Section 3) Appendix
B.2 Weakly Shill-Proof and Efficient Auctions (Section 3.2) Appendix
@@
In order to build towards a proof Theorem 3.6, we will prove that for a certain class of value distributions, every weakly shill-proof and efficient auction must have part of its game tree be a Dutch auction. Formally, we assume that the value distribution is sparse:
Definition B.3.
@@
$v^{k}-(v^{k+1}-v^{k})\frac{f(v^{k+1})}{f(v^{k})}<0.$ (3)
@@
Lemma B.4.
@@
$\mathds{E}\left[\sum_{i\in R}\tilde{t}_{i}(v)\mid V=\left(V_{-j},\left\{\gamma\right\}\right)\right]<\mathds{E}\left[\sum_{i\in R}\tilde{t}_{i}(v)\mid V=\left(V_{-j},\left\{\underline{W}\right\}\right)\right].$
Thus, the following shilling strategy is profitable compared to always reporting 00: if there exists $(V,\xi,W)$ such that $(W,\cdot)\in\mu(V,\xi)$ and $\underline{W}\in(0,\rho^{*})$, then select $W$. Otherwise, select the partition containing 00.
Proof.
Consider any $i\in R$ and $V$ and let $C=\left(\sum_{v^{k}\in V_{i}}f(v^{k})\right)^{-1}$. Then, applying Equation 2,
$\mathds{E}$ $\left[\tilde{t}_{i}(v)\mid V\right]+U_{i}(v^{j_{1}};V)=\mathds{E}\left[T_{i}(v_{i};V)\right]+U_{i}(v^{j_{1}};V)=C\sum_{m}f(v^{j_{m}})T_{i}(v^{j_{m}};V)$
$=C\sum_{m:v^{j_{m}}\in V_{i}}f(v^{j_{m}})\left(X_{i}(v^{j_{m}};V)v^{j_{m}}-\sum_{k<m}\left[\tilde{X}^{k}_{i}(V)\cdot(v^{j_{k+1}}-v^{j_{k}})\right]\right)$
$=C\left[\sum_{m:v^{j_{m}}\in V_{i}}f(v^{j_{m}})X_{i}(v^{j_{m}};V)v^{j_{m}}-\sum_{m:v^{j_{m}}\in V_{i}}\sum_{k<m}f(v^{j_{m}})\left[\tilde{X}^{k}_{i}(V)\cdot(v^{j_{k+1}}-v^{j_{k}})\right]\right]$
$=C\sum_{m:v^{j_{m}}\in V_{i}}\left[v^{j_{m}}X_{i}(v^{j_{m}};V)-(v^{j_{m+1}}-v^{j_{m}})\frac{F(\overline{{V}}_{i})-F(v^{j_{m}})}{f(v^{j_{m}})}\tilde{X}^{m}_{i}(V)\right]f(v^{j_{m}}).$
Applying the definition of the efficient allocation rule $\tilde{x}^{E}$, we know that for $(v^{m},i)\rhd(\gamma,j)$ and $(v^{m},i)\rhd(\gamma^{\prime},j)$, we can define $X_{i}(v^{m};V_{-j})\equiv X_{i}(v^{m};V_{-j},\left\{\gamma\right\})=X_{i}(v^{m};V_{-j},\left\{\gamma^{\prime}\right\}).$ Note that $\underline{W}\leq\min_{i}\left\{\overline{{V}}_{i}\right\}$ by assumption and therefore, for $\underline{W}\in\left(\gamma,\rho^{*}\right)$,
$\mathds{E}$ $\big{[}\tilde{t}_{i}(v)\mid V=\left(V_{-j},\left\{\gamma\right\}\right)\big{]}-\mathds{E}\left[\tilde{t}_{i}(v)\mid V=\left(V_{-j},\left\{\underline{W}\right\}\right)\right]$
$=C\sum_{m:\gamma\leq v^{j_{m}}<\underline{W}}\left[v^{j_{m}}X_{i}(v^{j_{m}};V)-(v^{j_{m+1}}-v^{j_{m}})\frac{F(\overline{{V}}_{i})-F(v^{j_{m}})}{f(v^{j_{m}})}\tilde{X}^{m}_{i}(V)\right]f(v^{j_{m}})$
$\leq C\sum_{m:\gamma\leq v^{j_{m}}<\underline{W}}\left[v^{j_{m}}-(v^{j_{m}+1}-v^{j_{m}})\frac{f(v^{j_{m}+1})}{f(v^{j_{m}})}\right]f(v^{j_{m}})X_{i}(v^{j_{m}};V_{-j})<0$
where the final inequality comes from sparsity. And so,
$\mathds{E}\left[\sum_{i\in R}\tilde{t}_{i}(v)\mid V=\left(V_{-j},\left\{\gamma\right\}\right)\right]<\mathds{E}\left[\sum_{i\in R}\tilde{t}_{i}(v)\mid V=\left(V_{-j},\left\{\underline{W}\right\}\right)\right],$
as claimed in the lemma. Thus, committing to misreport as $\underline{W}$ is strictly beneficial compared to any strategy that can only report $\gamma<\underline{W}$. ∎
Lemma B.5.
If $F$ is regular and sparse, then every public, weakly shill-proof, and efficient auction is a semi-Dutch auction with cutoff $\rho^{*}$.
Proof.
Suppose $F$ is regular and sparse. Consider an arbitrary weakly shill-proof and efficient auction, $(\tilde{x}^{E},\tilde{t},\mu,\xi_{0})$, and consider any $v$ such that $\max_{i}\left\{v_{i}\right\}<\rho^{*}$.
Proof of (i). First, we prove that, for any player $\xi$ and set of possible values $V$ such that there exists $(W,\cdot)\in\mu(V,\xi)$ where $0<\underline{W}<\rho^{*}$, it is the case that $V\subseteq\widecheck{V}$. Towards contradiction, suppose that there exists a $(\xi,V,W)$ such that $V\not\subseteq\widecheck{V}$, $\left(W,\cdot\right)\in\mu(V,\xi)$, and $\underline{W}\in\left(0,\rho^{*}\right)$. Let us suppose $\xi\in S$.
Let us first prove it is without loss to assume $(\xi,V,W)$ is such that for all $i$, $\underline{{V}}_{i}=0$ or $\underline{{V}}_{i}\geq\rho^{*}$. If there exists $(\xi,V,W)$ with $i$ such that $\underline{{V}}_{i}\in(0,\rho^{*})$, let us label that set as $V^{K}$ and let $V^{0}\supset V^{1}\supset\ldots\supset V^{K}$ be the sequence of on-path possible value sets to $V^{K}$. Let the players called along the path be $\xi^{0},\xi^{1},\ldots,\xi^{K}$ and the value partition selected by player $k$ to be $W^{k}$. Note that $V^{0}=\mathcal{V}^{N}$ is such that for all $i$, $\tilde{\underline{{V}}}_{i}=0$ or $\tilde{\underline{{V}}}_{i}\geq\rho^{*}$. So, the set $\mathcal{K}=\left\{k<K:\underline{W}^{k}\in(0,\rho^{*})\right\}\neq\emptyset$ and therefore $k^{*}=\min_{k\in\mathcal{K}}\left\{k\right\}$ is well-defined. If $k$ is such that $\underline{W}^{k}\notin(0,\rho^{*})$ and for all $i$, $\underline{{V}}^{k}_{i}\notin(0,\rho^{*})$, then it must be the case that for all $i$, $\underline{{V}}^{k+1}_{i}\notin(0,\rho^{*})$. Since $k^{*}$ is the first time in the game that a player selects a partition with $\underline{W}^{k}\in(0,\rho^{*})$, it must be the case that for all $i$, $\underline{{V}}^{k}_{i}=0$ or $\tilde{\underline{{V}}}_{i}\geq\rho^{*}$. Since $V^{k^{*}}\supset V^{K}$, $V^{k^{*}}\not\subseteq\widecheck{V}$. Thus, $(\xi^{k^{*}},V^{k^{*}},W^{k})$ is such that $V^{k^{*}}\not\subseteq\widecheck{V}$, $\left(W^{k^{*}},\cdot\right)\in\mu\left(V^{k^{*}},\xi^{k^{*}}\right)$, $\underline{W}^{k^{*}}\in\left(0,\rho^{*}\right)$, and for all $i$, $\underline{{V}}^{k^{*}}_{i}=0$ or $\underline{{V}}^{k^{*}}_{i}\geq\rho^{*}$.
So, in order to have $\underline{W}\in(0,\rho^{*})$, it must be the case that $0\in V_{\xi}$. Thus it is possible for $\xi$ to be a shill bidder while having so far only selected partitions that contain 00. Let $S=\left\{i:\overline{{V}}_{i}<\rho^{*}\right\}\cup\left\{\xi\right\}$. By assumption that $V\nsubseteq\widecheck{V}$, there must exist $i$ such that $\overline{{V}}_{i}\geq\rho^{*}$ and thus we can suppose that $R\neq\emptyset$. By assumption that $\underline{W}\in\left(0,\rho^{*}\right)$, we can suppose that $R$ is such that for all $i\in R$, $\overline{{V}}_{i}>\underline{W}$. By Lemma B.4, this would contradict the hypothesis that the auction is weakly shill-proof and so we must have $V\subseteq\widecheck{V}$ when there exists $(W,\cdot)\in\mu(V,\xi)$ such that $0<\underline{W}<\rho^{*}$.
Proof of (ii). We now prove that for any player $\xi$ and set of possible values $V\subseteq\widecheck{V}$, it is the case that $\mu(V,\xi)=\mu^{D}(V,\xi)$. Consider any option $\left(W,\cdot\right)\in\mu\left(V,\xi\right)$. Observe that by Lemma B.4, it is not the case that $0<\underline{W}<\overline{{V}}_{-\xi}$. So, $\underline{W}\geq\overline{{V}}_{-\xi}$. Since this is the case for all $V$, it must therefore be true that $\underline{W}=\overline{{V}}_{\xi}$. This is because if $\overline{{V}}_{\xi}>\underline{{V}}\geq\overline{{V}}_{-\xi}$, then there must have existed some earlier menu $\left(\tilde{W},\cdot\right)\in\mu\left(\tilde{V},\tilde{\xi}\right)$ for which $\underline{\tilde{W}}<\overline{\tilde{V}}_{\tilde{\xi}}$.

### v3

@@ Appendix B Strongly Shill-Proof Auctions (Section 3) Appendix
Appendix C Weakly Shill-Proof Auctions (Section 4) Appendix
C.1 Single-Action Auctions (Section 4.1) Appendix
Lemma C.1.
For any single-action auction, there exist unique $\tilde{x}:\vartheta^{N}\to\{0,1\}^{N}$ and $\tilde{t}:\vartheta^{N}\to\mathbb{R}^{N}$ such that:
  1. (i)
    (Correspondence) For all $v\in\vartheta^{N}$, $\tilde{x}^{*}(\theta)=x(\sigma(\theta;B))$ and $\tilde{t}(\theta)=t\left(\sigma(\theta;B)\right)$.
  2. (ii)
    (Individual Rationality) For all $i\in B$ and $\theta$, $\tilde{x}^{*}_{i}(\theta)\theta_{i}-\tilde{t}_{i}(\theta)\geq 0.$
  3. (iii)
    (Incentive Compatibility) For all $R$, $i\in R$, $\theta_{i}$, and $\theta_{i}^{\prime}$,
    $\mathds{E}_{\theta_{-i},\tilde{R}}\left[\tilde{x}^{*}_{i}\left(\sigma(\theta;\tilde{R})\right)\theta_{i}-\tilde{t}_{i}(\sigma(\theta;\tilde{R}))\mid\theta_{i},s_{i}=\psi(\theta_{j<i})\right]\\ \geq\mathds{E}_{\theta_{-i},\tilde{R}}\left[\tilde{x}^{*}_{i}\left(\sigma(\theta_{i}^{\prime},\theta_{-i};\tilde{R})\right)\theta_{i}-\tilde{t}_{i}\left(\sigma(\theta_{i}^{\prime},\theta_{-i};\tilde{R})\right)\mid\theta_{i},s_{i}=\psi(\theta_{j<i})\right].$ (3)
Proof.
[…]
Lemma C.2.
[…]
$\mathds{E}_{\theta}\left[\sum_{k\in R}\tilde{t}_{k}\left(\left\{\theta_{i}\right\}_{i\in S},\left\{\theta_{i}\right\}_{i\notin S}\right)\mid\theta_{j<\min S}\right]\leq\mathds{E}_{\theta}\left[\sum_{k\in R}\tilde{t}_{k}\left(0,\left\{\theta_{i}\right\}_{i\notin S}\right)\mid\theta_{j<\min S}\right].$
Proof.
Towards contradiction, suppose there exists $R$, $\theta_{j<\min S}$, and $\left\{\theta_{i}\right\}_{i\in S}$ such that
$\mathds{E}_{\theta}\left[\sum_{k\in R}\tilde{t}_{k}\left(\left\{\theta_{i}\right\}_{i\in S},\left\{\theta_{i}\right\}_{i\notin S}\right)\mid\theta_{j<\min S}\right]>\mathds{E}_{\theta}\left[\sum_{k\in R}\tilde{t}_{k}\left(0,\left\{\theta_{i}\right\}_{i\notin S}\right)\mid\theta_{j<\min S}\right].$
We now prove that the deviation by the coalition $S$ where they report $\left\{\theta_{i}\right\}_{i\in S}$ is profitable and therefore that the auction is not weakly shill-proof. By assumption, a shill bidder observes actions by all bidders who take actions before her. So, $\left\{\theta_{i}\right\}_{i\in S}$ can condition on $\theta_{j<\min S}$ when making decisions. Then, the strategy by $S$ of committing to report $\left\{\theta_{i}\right\}_{i\in S}$ regardless of what other bidders play after $\min_{i\in S}\left\{i\right\}$ must be strictly profitable compared to always reporting 0. Thus, we have found a strategy that does strictly better than always reporting 0: When the values before $\min_{i\in S}\left\{i\right\}$ are reported as $\theta_{j<\min S}$, report $\left\{\theta_{i}\right\}_{i\in S}$. Otherwise, report 0. This strategy in the direct game immediately translates to a profitable deviation in the auction by Definition A.1 and Lemma C.1 and thus the equilibrium is not weakly shill-proof. ∎
Now, when discussing single-action auctions, we focus on the direct mechanisms associated to weakly shill-proof auctions and so we will refer to an auction as $(\tilde{x},\tilde{t},\psi)$ without reference to $R$.
Lemma C.3.
Suppose a single-action, orderly auction $(\tilde{x},\tilde{t},\psi)$ is weakly shill-proof. Then, for all $i$, $\theta$, and $\theta_{j>i}^{\prime}$, $\left[\tilde{x}_{i}\left(\theta\right)=\tilde{x}_{i}\left(\theta_{j\leq i},\theta_{j>i}^{\prime}\right)\implies\tilde{t}_{i}\left(\theta\right)=\tilde{t}_{i}\left(\theta_{j\leq i},\theta_{j>i}^{\prime}\right)\right].$
Proof.
Towards contradiction, suppose there exists $i$, $\theta$, and $\theta_{j>i}^{\prime}$, such that $\tilde{x}_{i}\left(\theta\right)=\tilde{x}_{i}\left(\theta_{j\leq i},\theta_{j>i}^{\prime}\right)$, but $\tilde{t}_{i}\left(\theta;s\right)>\tilde{t}_{i}\left(\theta_{j\leq i},\theta_{j>i}^{\prime};s\right)$. Because the auction is orderly, we can apply Lemma B.1 to conclude that $\tilde{x}_{i}\left(\theta\right)=\tilde{x}_{i}\left(\theta_{j\leq i},\theta_{j>i}^{\prime}\right)=1$. Let $R=\left\{1,\ldots,i\right\}$. Then,
$\mathds{E}_{\theta}\left[\sum_{k\in R}\tilde{t}_{k}(\left\{\theta_{i}\right\}_{i\in S},\left\{\theta_{i}\right\}_{i\notin S})\mid\theta_{j\leq\min S}\right]$ $=\tilde{t}_{i}\left(\theta\right)>\tilde{t}_{i}\left(\theta_{j\leq i},\theta_{j>i}^{\prime}\right)\geq\tilde{t}_{i}\left(\theta_{j\leq i},0\right).$
This violates Lemma C.2, and so we have reached a contradiction. ∎
Lemma C.4.
Suppose a single-action, orderly auction $(\tilde{x},\tilde{t},\psi)$ is mildly ex-post incentive compatible and weakly shill-proof. Then, there exists $i<N$ such that for all $\theta_{i}$, $\theta_{i}^{\prime}$, and $\theta_{-i}\in\psi^{-1}_{i}(\theta_{j<i})$, $\left[\tilde{x}_{i}(\theta)=\tilde{x}_{i}(\theta_{i}^{\prime},\theta_{-i})\implies\tilde{t}_{i}(\theta)=\tilde{t}_{i}(\theta_{i}^{\prime},\theta_{-i})\right].$
Proof.
Consider $i<N,R\ni i,\theta_{i},\theta_{i}^{\prime}$, and $\theta_{-i}\in\psi^{-1}_{i}(\theta_{j<i})$ such that $\tilde{x}_{i}(\theta)=\tilde{x}_{i}(\theta_{i}^{\prime},\theta_{-i})$. WLOG, suppose $\theta_{i}>\theta_{i}^{\prime}$. By monotonicity, $\tilde{t}_{i}(\theta)\geq\tilde{t}_{i}(\theta_{i}^{\prime},\theta_{-i})$. Towards contradiction, suppose $\tilde{t}_{i}(\theta)>\tilde{t}_{i}(\theta_{i}^{\prime},\theta_{-i})$. By the winner-paying property, $\tilde{t}_{i}(\theta)>\tilde{t}_{i}(\theta_{i}^{\prime},\theta_{-i})$ implies that $\tilde{x}_{i}(\theta)=\tilde{x}_{i}(\theta_{i}^{\prime},\theta_{-i})=1$. However, note that $\tilde{t}_{i}(\theta)>\tilde{t}_{i}(\theta_{i}^{\prime},\theta_{-i})$ would mean that the utility of reporting $\theta_{i}^{\prime}$ would be higher than truthful reporting under true value $\theta_{i}$ which would violate the mildly ex-post incentive compatiblility and thus $\tilde{t}_{i}(\theta_{i}^{\prime},\theta_{-i})=\tilde{t}_{i}(\theta_{i}^{\prime},\theta_{-i})$. ∎
Proof of Theorem 4.3.
Towards contradiction, suppose such an auction did exist. Fix $i<N$ and $\theta_{j<i}$ and then suppose that $\theta_{i}<\theta^{M}$. Combining Lemmatta C.3 and C.4, we can see that for all $\theta_{i}^{\prime}$ and $\theta_{-i},\theta_{-i}^{\prime}\in\psi^{-1}_{i}(\theta_{j<i})$, if $\tilde{x}_{i}(\theta)=\tilde{x}_{i}(\theta^{\prime})$, then $\tilde{t}_{i}(\theta)=\tilde{t}_{i}(\theta^{\prime})$. So, define $\tilde{t}_{i}$ as the (constant) $\tilde{t}_{i}(\theta)$ for all $\theta$ such that $\tilde{x}_{i}(\theta)=1$.
In order for bidder $i$ to have an ex-post strategy, when $\tilde{x}(\theta)=1$, it must also be the case that $\tilde{x}_{i}(\theta^{M},\theta_{-i})=1$. So, applying the winner-paying property (and suppressing that the expectation is conditioned on $s_{i}=\psi_{i}(\theta_{j<i})$), we have
$\mathds{E}_{\theta_{-i}}\left[\tilde{x}_{i}\left(\theta^{M},\theta_{-i}\right)\theta_{i}-\tilde{t}_{i}\left(\theta^{M},\theta_{-i}\right)\right]\\ =\mathds{E}_{\theta_{-i}}\left[\theta_{i}-\tilde{t}_{i}^{*}\mid\tilde{x}_{i}(\theta)=1\right]+\mathds{E}_{\theta_{-i}}\left[\theta_{i}-\tilde{t}_{i}^{*}\mid\tilde{x}_{i}(\theta)=0,\tilde{x}_{i}\left(\theta^{M},\theta_{-i}\right)=1\right],$ (4)
$\text{and }\mathds{E}_{\theta_{-i}}\left[\tilde{x}_{i}\left(\theta\right)\theta_{i}-\tilde{t}_{i}\left(\theta\right)\right]=\mathds{E}_{\theta_{-i}}\left[\theta_{i}-\tilde{t}_{i}^{*}\mid\tilde{x}_{i}(\theta)=1\right].$ (5)
Taking the difference between Equation 4 and Equation 5, we see that
$\mathds{E}_{\theta_{-i}}\left[\tilde{x}_{i}\left(\theta^{M},\theta_{-i}\right)\theta_{i}-\tilde{t}_{i}\left(\theta^{M},\theta_{-i}\right)\right]-\mathds{E}_{\theta_{-i}}\left[\tilde{x}_{i}\left(\theta\right)\theta_{i}-\tilde{t}_{i}\left(\theta\right)\right]\\ =\mathds{E}_{\theta_{-i}}\left[\theta_{i}-\tilde{t}_{i}^{*}\mid\tilde{x}_{i}(\theta)=0,\tilde{x}_{i}\left(\theta^{M},\theta_{-i}\right)=1\right]$ (6)
Now, by definition of orderly, $\tilde{x}$ is monotone, and by assumption $\tilde{t}$ is monotone. If there exists $\theta^{m}$ such that $\mathds{P}[\tilde{x}_{i}(\theta^{m},\theta_{-i})=1]<\mathds{P}[\tilde{x}_{i}(\theta^{M},\theta_{-i})=1]$, then for such $m$,
$\mathds{E}_{\theta_{-i}}\left[\tilde{x}_{i}(\theta^{m},\theta_{-i})\theta^{m}-\tilde{t}_{i}(\theta^{m},\theta_{-i})\right]\geq\mathds{E}_{\theta_{-i}}\left[\tilde{x}_{i}(\theta^{m-1},\theta_{-i})\theta^{m}-\tilde{t}_{i}(\theta^{m-1},\theta_{-i})\right]$
$>\mathds{E}_{\theta_{-i}}\left[\tilde{x}_{i}(\theta^{m-1},\theta_{-i})\theta^{m-1}-\tilde{t}_{i}(\theta^{m-1},\theta_{-i})\right]\geq 0,$
where the last inequality comes from the IR condition. Thus, the IR constraint does not bind for $\theta_{i}=\theta^{m}$. Since the good has to be allocated to the highest type, for all $i<N$, there exists $\theta_{-i}$ such that $\tilde{x}_{i}(\theta)=0$ and $\tilde{x}_{i}(\theta^{M},\theta_{-i})=1$. Thus,
$\mathds{E}_{\theta_{-i}}\left[\theta_{i}-\tilde{t}_{i}^{*}\mid\tilde{x}_{i}(\theta)=0,\tilde{x}_{i}\left(\theta^{M},\theta_{-i}\right)=1\right]>0.$ (7)
Combining Equations 6 and 7, we see that
$\mathds{E}_{\theta_{-i}}\left[\tilde{x}_{i}\left(\theta^{M},\theta_{-i}\right)\theta_{i}-\tilde{t}_{i}\left(\theta^{M},\theta_{-i}\right)\right]>\mathds{E}_{\theta_{-i}}\left[\tilde{x}_{i}\left(\theta\right)\theta_{i}-\tilde{t}_{i}\left(\theta\right)\right].$ (8)
$\mathds{E}_{\theta_{-i}}\left[\tilde{x}_{i}\left(\theta\right)\theta_{i}-\tilde{t}_{i}(\theta)\right]\geq\mathds{E}_{\theta_{-i}}\left[\tilde{x}_{i}\left(\theta_{i}^{\prime},\theta_{-i}\right)\theta_{i}-\tilde{t}_{i}\left(\theta_{i}^{\prime},\theta_{-i}\right)\right].$
Taking $\theta_{i}^{\prime}=\theta^{M}$, Equation 8 violates the IC constraint from Lemma C.1—and thus we have reached a contradiction. ∎
C.2 Efficient Auctions (Section 4.2.1) Appendix
@@ Appendix C Weakly Shill-Proof Auctions (Section 4) Appendix
In order to build towards a proof Proposition 4.6, we will prove that for a certain class of value distributions, every weakly shill-proof and efficient auction must be a semi-Dutch auction. Formally, we assume that the value distribution is sparse:
Definition C.5.
@@
$\theta^{k}-(\theta^{k+1}-\theta^{k})\frac{f(\theta^{k+1})}{f(\theta^{k})}<0.$ (9)
@@
Lemma C.6.
Consider an efficient auction and suppose $F$ is regular and sparse. Let $R$, $\Theta$ such that $\Theta_{i}=\left\{w:w\in[\underline{{\Theta}}_{i},\overline{{\Theta}}_{i}]\right\}$ for all $i\in R$, and consider $(\underline{W},j)$ such that $\underline{W}<\theta^{\rho^{*}},j\notin R$ and for all $i\in R,(\overline{{\Theta}}_{i},i)\rhd(\underline{W},j)$. Then, for all $\gamma<\underline{W}$,
$\mathds{E}\left[\sum_{i\in R}\tilde{t}_{i}(\theta)\mid\Theta=\left(\Theta_{-j},\left\{\gamma\right\}\right)\right]<\mathds{E}\left[\sum_{i\in R}\tilde{t}_{i}(\theta)\mid\Theta=\left(\Theta_{-j},\left\{\underline{W}\right\}\right)\right].$
Thus, the following shilling strategy is profitable compared to always reporting 0: if there exists $(\Theta,\xi,W)$ such that $(W,\cdot)\in\mu(\Theta,\xi)$ and $\underline{W}\in(0,\theta^{\rho^{*}})$, then select $W$. Otherwise, select the partition containing 0.
Proof.
Consider any $i\in R$ and $\Theta$ and let $C=\left(\sum_{\theta^{k}\in\Theta_{i}}f(\theta^{k})\right)^{-1}$. Then, applying Equation 2,
$\mathds{E}$ $\left[\tilde{t}_{i}(\theta)\mid\Theta\right]+U_{i}(\theta^{j_{1}};\Theta)=\mathds{E}\left[T_{i}(\theta_{i};\Theta)\right]+U_{i}(\theta^{j_{1}};\Theta)=C\sum_{m}f(\theta^{j_{m}})T_{i}(\theta^{j_{m}};\Theta)$
$=C\sum_{m:\theta^{j_{m}}\in\Theta_{i}}f(\theta^{j_{m}})\left(X_{i}(\theta^{j_{m}};\Theta)\theta^{j_{m}}-\sum_{k<m}\left[\tilde{X}^{k}_{i}(\Theta)\cdot(\theta^{j_{k+1}}-\theta^{j_{k}})\right]\right)$
$=C\left[\sum_{m:\theta^{j_{m}}\in\Theta_{i}}f(\theta^{j_{m}})X_{i}(\theta^{j_{m}};\Theta)\theta^{j_{m}}-\sum_{m:\theta^{j_{m}}\in\Theta_{i}}\sum_{k<m}f(\theta^{j_{m}})\left[\tilde{X}^{k}_{i}(\Theta)\cdot(\theta^{j_{k+1}}-\theta^{j_{k}})\right]\right]$
$=C\sum_{m:\theta^{j_{m}}\in\Theta_{i}}\left[\theta^{j_{m}}X_{i}(\theta^{j_{m}};\Theta)-(\theta^{j_{m+1}}-\theta^{j_{m}})\frac{F(\overline{{\Theta}}_{i})-F(\theta^{j_{m}})}{f(\theta^{j_{m}})}\tilde{X}^{m}_{i}(\Theta)\right]f(\theta^{j_{m}}).$
Applying the definition of the efficient allocation rule $\tilde{x}^{E}$, we know that for $(\theta^{m},i)\rhd(\gamma,j)$ and $(\theta^{m},i)\rhd(\gamma^{\prime},j)$, we can define $X_{i}(\theta^{m};\Theta_{-j})\equiv X_{i}(\theta^{m};\Theta_{-j},\left\{\gamma\right\})=X_{i}(\theta^{m};\Theta_{-j},\left\{\gamma^{\prime}\right\}).$ Note that $\underline{W}\leq\min_{i}\left\{\overline{{\Theta}}_{i}\right\}$ by assumption and therefore, for $\underline{W}\in\left(\gamma,\theta^{\rho^{*}}\right)$,
$\mathds{E}$ $\big[\tilde{t}_{i}(\theta)\mid\Theta=\left(\Theta_{-j},\left\{\gamma\right\}\right)\big]-\mathds{E}\left[\tilde{t}_{i}(\theta)\mid\Theta=\left(\Theta_{-j},\left\{\underline{W}\right\}\right)\right]$
$=C\sum_{m:\gamma\leq\theta^{j_{m}}<\underline{W}}\left[\theta^{j_{m}}X_{i}(\theta^{j_{m}};\Theta)-(\theta^{j_{m+1}}-\theta^{j_{m}})\frac{F(\overline{{\Theta}}_{i})-F(\theta^{j_{m}})}{f(\theta^{j_{m}})}\tilde{X}^{m}_{i}(\Theta)\right]f(\theta^{j_{m}})$
$\leq C\sum_{m:\gamma\leq\theta^{j_{m}}<\underline{W}}\left[\theta^{j_{m}}-(\theta^{j_{m}+1}-\theta^{j_{m}})\frac{f(\theta^{j_{m}+1})}{f(\theta^{j_{m}})}\right]f(\theta^{j_{m}})X_{i}(\theta^{j_{m}};\Theta_{-j})<0$
where the final inequality comes from sparsity. And so,
$\mathds{E}\left[\sum_{i\in R}\tilde{t}_{i}(\theta)\mid\Theta=\left(\Theta_{-j},\left\{\gamma\right\}\right)\right]<\mathds{E}\left[\sum_{i\in R}\tilde{t}_{i}(\theta)\mid\Theta=\left(\Theta_{-j},\left\{\underline{W}\right\}\right)\right],$
as claimed in the statement of the lemma. Thus, committing to misreport as $\underline{W}$ is strictly beneficial compared to any strategy that can only report $\gamma<\underline{W}$. ∎
Lemma C.7.
If $F$ is regular and sparse, then every public, weakly shill-proof, and efficient auction is a semi-Dutch auction with cutoff $\theta^{\rho^{*}}$.
Proof.
Suppose $F$ is regular and sparse. Consider an arbitrary weakly shill-proof and efficient auction, $(\tilde{x}^{E},\tilde{t},\mu,\xi_{0})$, and consider any $\theta$ such that $\max_{i}\left\{\theta_{i}\right\}<\theta^{\rho^{*}}$. We prove that both parts of the definition of a semi-Dutch auction are necessary.
Proof of Condition i of Definition 4.5. First we prove that if, for any player $\xi$ and set of possible values $\Theta$, there exists $(W,\cdot)\in\mu(\Theta,\xi)$ where $0<\underline{W}<\theta^{\rho^{*}}$, then $\Theta\subseteq\widecheck{\Theta}$. Towards contradiction, suppose that there exists a $(\xi,\Theta,W)$ such that $\Theta\not\subseteq\widecheck{\Theta}$, $\left(W,\cdot\right)\in\mu(\Theta,\xi)$ $\underline{W}\in\left(0,\theta^{\rho^{*}}\right)$, and $\xi\in S$.
Let us first prove it is without loss to assume $(\xi,\Theta,W)$ is such that for all $i$, $\underline{{\Theta}}_{i}=0$ or $\underline{{\Theta}}_{i}\geq\theta^{\rho^{*}}$. If there exists $(\xi,\Theta,W)$ and ii such that $\underline{{\Theta}}_{i}\in(0,\theta^{\rho^{*}})$, let us label that set as $\Theta^{K}$ and let $\Theta^{0}\supset\Theta^{1}\supset\ldots\supset\Theta^{K}$ be the sequence of on-path possible value sets preceding $\Theta^{K}$. Let the players called along the path be $\xi^{0},\xi^{1},\ldots,\xi^{K}$ and the value partition selected by player $k$ to be $W^{k}$. Note that $\Theta^{0}=\vartheta^{N}$ and so for all $i$, $\tilde{\underline{{\Theta}}}_{i}=0$ or $\tilde{\underline{{\Theta}}}_{i}\geq\theta^{\rho^{*}}$. So, the set $\mathcal{K}=\left\{k<K:\underline{W}^{k}\in(0,\theta^{\rho^{*}})\right\}\neq\emptyset$ and therefore $k^{*}=\min_{k\in\mathcal{K}}\left\{k\right\}$ is well-defined. If $k$ is such that $\underline{W}^{k}\notin(0,\theta^{\rho^{*}})$ and for all $i$, $\underline{{\Theta}}^{k}_{i}\notin(0,\theta^{\rho^{*}})$, then it must be the case that for all $i$, $\underline{{\Theta}}^{k+1}_{i}\notin(0,\theta^{\rho^{*}})$. Since $k^{*}$ is the first time in the game that a player selects a partition with $\underline{W}^{k}\in(0,\theta^{\rho^{*}})$, it must be the case that for all $i$, $\underline{{\Theta}}^{k}_{i}=0$ or $\tilde{\underline{{\Theta}}}_{i}\geq\theta^{\rho^{*}}$. Since $\Theta^{k^{*}}\supset\Theta^{K}$, $\Theta^{k^{*}}\not\subseteq\widecheck{\Theta}$. Thus, $(\xi^{k^{*}},\Theta^{k^{*}},W^{k})$ is such that $\Theta^{k^{*}}\not\subseteq\widecheck{\Theta}$, $\left(W^{k^{*}},\cdot\right)\in\mu\left(\Theta^{k^{*}},\xi^{k^{*}}\right)$, $\underline{W}^{k^{*}}\in\left(0,\theta^{\rho^{*}}\right)$, and for all $i$, $\underline{{\Theta}}^{k^{*}}_{i}=0$ or $\underline{{\Theta}}^{k^{*}}_{i}\geq\theta^{\rho^{*}}$.
So, in order to have $\underline{W}\in(0,\theta^{\rho^{*}})$, it must be the case that $0\in\Theta_{\xi}$. Thus it is possible for $\xi$ to be a shill bidder while having so far only selected partitions that contain 0. Let $S=\left\{i:\overline{{\Theta}}_{i}<\theta^{\rho^{*}}\right\}\cup\left\{\xi\right\}$. By assumption that $\Theta\nsubseteq\widecheck{\Theta}$, there must exist $i$ such that $\overline{{\Theta}}_{i}\geq\theta^{\rho^{*}}$ and thus we can suppose that $R\neq\emptyset$. By assumption that $\underline{W}\in\left(0,\theta^{\rho^{*}}\right)$, we can suppose that $R$ is such that for all $i\in R$, $\overline{{\Theta}}_{i}>\underline{W}$. By Lemma C.6, this would contradict the hypothesis that the auction is weakly shill-proof and so we must have $\Theta\subseteq\widecheck{\Theta}$ when there exists $(W,\cdot)\in\mu(\Theta,\xi)$ such that $0<\underline{W}<\theta^{\rho^{*}}$.